OpenStudy (anonymous):

Which equation is a quadratic equation? A.y = 2x(5 - x) + 7x B.y = 8x + 2 C.y + x2 = (x - 2)(x + 8) D.y - 7x = (x2 + 1)(x - 3)

jimthompson5910 (jim_thompson5910):

which one do you think it is?

OpenStudy (anonymous):

C? O.o

OpenStudy (anonymous):

A quadratic equation must have 2 as its highest degree. :)

jimthompson5910 (jim_thompson5910):

no, here's why y + x^2 = (x - 2)(x + 8) y + x^2 = x^2 + 8x - 2x - 16 y = x^2 + 8x - 2x - 16 - x^2 y = 6x - 16 .... which is a linear (and NOT a quadratic) equation

OpenStudy (anonymous):

B is not (highest power of x is 1.) C is not (x^2 cancels out.) D is not (x^3 is present.)

OpenStudy (anonymous):

A? (:

jimthompson5910 (jim_thompson5910):

yes, choice A is a quadratic equation